# GestureLSM: Latent Shortcut based Co-Speech Gesture Generation with Spatial-Temporal Modeling

### v2

@@ 2 Related Works
Existing works on co-speech gesture generation mostly employ skeleton- or joint-level pose representations. [14] use an adversarial framework to predict hand and arm poses from audio, and leverage conditional generation [3] based on pix2pixHD [48] for videos. Some recent works [31, 9, 50, 28, 55, 29, 39] learn the hierarchical semantics or leverage contrastive learning to obtain joint audio-gesture embeding to assist the gesture pose generation. HA2G [31] construct high and low level audio-motion embedding for gesture generation. TalkShow [52] estimates SMPL [35] poses, and models the body and hand motions for talk-show videos. CaMN [26] and EMAGE [27] propose a large scale conversational and speech datasets for joint face and body modeling with diverse style control, with GPT-style decoding for gesture generation.
@@
Diffusion models [17] have demonstrated impressive generation quality across various modalities [36, 1], but suffer from slow inference speeds due to their iterative sampling process. To address these efficiency challenges, several approaches have been proposed, such as Consistency Models [44, 34] and Diffusion Distillations [37, 53]. However, these methods still face limitations in terms of training speed and flexibility.
One promising direction for improving inference speed is the use of flow matching. [30] introduces rectified flow, which provides a crucial baseline for diffusion acceleration. Through rectification, they straighten the ODE path of flow-matching-based diffusion models. Building on this, [32, 57, 47] scales rectified flow to large-scale text-to-image generation, achieving one-step generation.
@@ 3.3 Gesture Latent Shortcut Model
where $\alpha$, $\beta$ controls the skewness of the sampling. We compared different sampling methods and the skewness effect for model learning. From Tab.2(c) and Tab.3(e), we discover that left skewness with emphasis when t approaches 1 can significantly improve the generation quality.
@@ 4.2 Quantitative Comparisons
We evaluate the realism of generated body gestures using the Fréchet Gesture Distance (FGD)[54], which quantifies the distributional similarity between ground truth and synthesized gestures. Diversity (Div.)[22] is measured by calculating the average L1 distance across multiple gesture clips. To assess speech-motion synchronization, we use Beat Constancy (BC) [23]. For facial motion evaluation, we compute the vertex Mean Squared Error (MSE) [49] to assess positional accuracy. Additionally, the efficiency of our approach is quantified through the Average Inference Time per Sentence (AITS). While previous works consider Beat Constancy higher the better, we discover the significant bias with this metric and propose to compare this metric for the generated data with that of the ground-truth data. We defer the additional details in Appendix E.
@@ 4.4 Ablation Studies
To verify our designs, we conduct ablations. we defer additional ablation studies in the Appendix Sec.D.
@@ Appendix B Implementation Details
In the construction of the RVQVAEs, the codebook is initialized uniformly, with each entry having a feature length of 128 and a total size of 1,024 per body region. The codebook updates occur solely during the quantization process, with resets following the Momask strategy [15]. The RVQVAEs are trained for 30,000 iterations, with a learning rate of $2\times 10^{-4}$. The GestureLSM model contains 3 layers of cross-attention for audio-gesture feature fusion and 8 layers of spatial-temporal attention blocks. The latent dimension is set to 256 with feed-forward size of 1024. During the second training stage for speech-to-gesture generation, the codebook remains frozen. We train the GestureLSM model for 1000 epochs. We utilize the Adam optimizer with a learning rate of $2\times 10^{-4}$. All experiments are conducted on a single NVIDIA A100 GPU. We adopt a guidance dropout rate of 0.1 during training and a speech-conditioning ratio of 2 during generation.
@@ Appendix D Additional Experimental Results
We analyze the sequential design of the proposed two types od attentions. . Tab.3(d) shows attention in spatial-temporal order present slightly improvement.

### v3

@@ 2 Related Works
Existing works on co-speech gesture generation mostly employ skeleton- or joint-level pose representations. [15] use an adversarial framework to predict hand and arm poses from audio, and leverage conditional generation [3] based on pix2pixHD [56] for videos. Some recent works [35, 9, 58, 30, 63, 33, 47, 32, 31] learn the hierarchical semantics or leverage contrastive learning to obtain joint audio-gesture embeding to assist the gesture pose generation. HA2G [35] construct high and low level audio-motion embedding for gesture generation. TalkShow [60] estimates SMPL [43] poses, and models the body and hand motions for talk-show videos. CaMN [28] and EMAGE [29] propose a large scale conversational and speech datasets for joint face and body modeling with diverse style control, with GPT-style decoding for gesture generation.
@@
Diffusion models [18, 38, 24, 39, 40] have demonstrated impressive generation quality across various modalities [44, 1, 42], but suffer from slow inference speeds due to their iterative sampling process. To address these efficiency challenges, several approaches have been proposed, such as Consistency Models [52, 41] and Diffusion Distillations [45, 61]. However, these methods still face limitations in terms of training speed and flexibility.
One promising direction for improving inference speed is the use of flow matching. [34] introduces rectified flow, which provides a crucial baseline for diffusion acceleration. Through rectification, they straighten the ODE path of flow-matching-based diffusion models. Building on this, [36, 65, 55, 14, 66] scales rectified flow to large-scale text-to-image generation, achieving one-step generation.
@@ 3.3 Gesture Latent Shortcut Model
where $\alpha$, $\beta$ controls the skewness of the sampling. We compared different sampling methods and the skewness effect for model learning. From Tab.2(c), we discover that left skewness with emphasis when t approaches 1 can significantly improve the generation quality.
@@ 4.2 Quantitative Comparisons
We evaluate the realism of generated body gestures using the Fréchet Gesture Distance (FGD)[62], which quantifies the distributional similarity between ground truth and synthesized gestures. Diversity (Div.)[23] is measured by calculating the average L1 distance across multiple gesture clips. To assess speech-motion synchronization, we use Beat Constancy (BC) [25]. For facial motion evaluation, we compute the vertex Mean Squared Error (MSE) [57] to assess positional accuracy. Additionally, the efficiency of our approach is quantified through the Average Inference Time per Sentence (AITS). While previous works consider Beat Constancy higher the better, we discover the significant bias with this metric and propose to compare this metric for the generated data with that of the ground-truth data. We defer the additional details in the Appendix.
@@ 4.4 Ablation Studies
To verify our designs, we conduct ablations. we defer additional ablation studies in the Appendix.
@@ Appendix B Implementation Details
In the construction of the RVQVAEs, the codebook is initialized uniformly, with each entry having a feature length of 128 and a total size of 1,024 per body region. The codebook updates occur solely during the quantization process, with resets following Contextual Gesture [33]. The RVQVAEs are trained for 30,000 iterations, with a learning rate of $2\times 10^{-4}$. The GestureLSM model contains 3 layers of cross-attention for audio-gesture feature fusion and 8 layers of spatial-temporal attention blocks. The latent dimension is set to 256 with feed-forward size of 1024. During the second training stage for speech-to-gesture generation, the codebook remains frozen. We train the GestureLSM model for 1000 epochs. We utilize the Adam optimizer with a learning rate of $2\times 10^{-4}$. All experiments are conducted on a single NVIDIA A100 GPU. We adopt a guidance dropout rate of 0.1 during training and a speech-conditioning ratio of 2 during generation.
@@ Appendix D Additional Experimental Results
We analyze the sequential design of the proposed two types od attentions. Tab.3(d) shows attention in spatial-temporal order present slightly improvement.